OpenStudy (anonymous):

what is the solution.. medal. 3x + 2y + z = 7 5x + 5y + 4z = 3 3x + 2y + 3z = 1

OpenStudy (imstuck):

Do you know how to solve 3 equations at a time? You must be in Algebra 2?

OpenStudy (anonymous):

i am in alg. 2 but idk how to do it :/

OpenStudy (imstuck):

Give me a couple of minutes to work it out then I will be able to take you step by step. Is that ok with You?

OpenStudy (anonymous):

yeah :) thanks!

OpenStudy (mertsj):

Multiply the first equation by -4 and post the result.

OpenStudy (anonymous):

multiply where though?

OpenStudy (imstuck):

He means the whole equation. Multiply everything in the first eqaution by a -4.

OpenStudy (mertsj):

The first equation which is 3x+2y+z=7

OpenStudy (anonymous):

-12x + -8y + (what you do with "z") = ?

OpenStudy (mertsj):

\[-4(3x+2x+z=7)\]=-12x-8x-4x=-28

OpenStudy (mertsj):

typo: -12x-8y-4z=-28

OpenStudy (mertsj):

Now write the second equation underneath that and add the two equations

OpenStudy (mertsj):



OpenStudy (mertsj):

Add those

OpenStudy (anonymous):

-7x + -3y + 0z = 31

OpenStudy (mertsj):

-28+3 is NOT 31

OpenStudy (anonymous):

-25 didnt see negative. im sorry.

OpenStudy (mertsj):

What is 0 times z?

OpenStudy (anonymous):

anything times 0 is 0... so 0?

OpenStudy (mertsj):

So now we have the equation: -7x-3y=- 25

OpenStudy (mertsj):

Do you notice that we have eliminated the z ?

OpenStudy (anonymous):

yeah

OpenStudy (mertsj):

We have to do that again and we have to use the third equation which hasn't yet been a part of the solution. So this time multiply the first equation by -3 and post the result.

OpenStudy (anonymous):

the original first equation?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

-9x + -6y + (again, forgot what you do with z) = -21

OpenStudy (mertsj):

z means 1z. Everyone agrees that if the coefficient is a 1 we won't bother to write it and everyone will just know that it is there.

OpenStudy (mertsj):

-3(3x+2y+1z=7) -9x-6y-3z=-21

OpenStudy (mertsj):

Now add the third equation to that result.

OpenStudy (anonymous):

-6x + -4y + 0z = -20

OpenStudy (mertsj):

yes. Do you notice that the z has been eliminated again since 0z = 0 ?

OpenStudy (anonymous):

yeah

OpenStudy (mertsj):

So now we have two equations with only x and y and they are: -7x - 3y = -25 and -6x - 4y = -20

OpenStudy (mertsj):

Can you solve that system?

OpenStudy (anonymous):

yeah... got it from here! thanks :)

OpenStudy (mertsj):

yw